# 2]organization=ASML, city=Veldhoven, country=The Netherlands

### v1

@@ 1 Introduction
Increasingly stringent performance requirements of next-generation industrial systems require data-driven models to integrate prior knowledge via interpretable physical principles. Embedding such principles improves model accuracy, increases interpretability, and typically reduces model complexity Soderstrom2001SystemIdentification . These characteristics are crucial not only for advanced multivariable control design methods VanDeWal2002MultivariableMotion ; Oomen2015InferentialInterest ; VanHerpen2014ExploitingControl but also for mechanical system design, model validation, machine monitoring techniques, and simulation studies Schmidt2011TheMechatronics ; Heertjes2020ControlDevelopments ; Steinbuch2022MotionLaw ; Classens2023FromReconfiguration .
@@
Despite extensive work on modal identification, relatively few methods explicitly enforce the rank constraint inherent to modal decompositions. Traditional approaches include single-step methods, which approximate the original nonlinear least-squares problem by a linear surrogate Guillaume2003AEstimator ; Peeters2004TheEstimation ; Peeters2005PolyMax:Analysis . Improved estimation accuracy has been achieved through iterative multi-step methods that directly minimize the nonlinear cost function, using gradient-descent optimization Bayard1994High-orderResults , pseudo-linear regression techniques Sanathanan1961TransferPolynomials ; Blom2010MultivariableRegression or vector fitting procedures Semlyen1999RationalFitting ; Gustavsen2004ADomain ; Civera2021ExperimentalMethod . These approaches have been applied in a range of multi-input multi-output (MIMO) modal identification methods Guillayme2000Frequency-domainAnalysis ; Verboven2005MultivariableEstimation ; Vayssettes2014Frequency-domainTests ; Pintelon2007UncertaintyAnalysis ; Shirvani2023LinearData ; Reynders2012SystemComparison . However, single-step methods generally lack statistical optimality guarantees, while many multi-step approaches either neglect the rank constraint or rely on singular value decompositions, which are typically effective only for systems with sufficiently spaced modes.
@@
Although many parametric modal identification techniques are available, few provide a systematic and computationally efficient framework that explicitly accounts for the rank condition and exploits the additive structure. This paper introduces a novel two-stage structured identification strategy for estimating MIMO input-output models in modal form. In the first stage, an additive model is estimated using a refined instrumental variable method Young1980RefinedAnalysis , which is subsequently projected onto a modal form using the indirect prediction error method (IPEM) Soderstrom1991AnIdentification . The framework accommodates both proportional damping and general viscous damping models DeKraker2004ADynamics , as these formulations are commonly encountered in industrial applications Verbaan2015RobustStages . The main contributions of this paper are:
@@ 1. C1:
    A frequency-domain refined instrumental variable method is derived for estimating continuous-time MIMO models in additive transfer function form.
@@ 2. C2:
    A two-stage structured identification framework is introduced that projects additive models onto the modal form, accommodating both proportional and general viscous damping models.
  3. C3:
    Experimental validation of the developed parametric modal system identification method on an experimental wafer-stage system.
@@ 3. C3:
The remainder of this paper is organized as follows. Section 2 introduces the experimental setup, presents the modal modeling framework for flexible mechanical systems, and provides a concise problem formulation. Section 3 formulates the two‐stage identification problem, with the solution procedures for the first stage provided in Section 4 and for the second stage in Section 5. Experimental validation is described in Section 6, and finally, Section 7 presents the conclusions.
Notation: Scalars, vectors, matrices and sets are written as $a$, $\mathbf{a}$, $\mathbf{A}$ and $\mathcal{A}$, respectively. The imaginary unit is denoted by $j$ and satisfies $j^{2}=-1$. For $\mathbf{z}\in\mathbb{C}^{n}$, the operation $\Re\{\mathbf{z}\}$ returns the real part, $\Im\{\mathbf{z}\}$ returns the imaginary part and $\bar{\mathbf{z}}$ returns the complex conjugate of the complex vector $\mathbf{z}$. For a matrix $\mathbf{A}$, its transpose is written as $\mathbf{A}^{\top}$, and its Hermitian (conjugate transpose) as $\mathbf{A}^{\mathsf{H}}$. If $\mathbf{x}\in\mathbb{C}^{n}$ and $\mathbf{Q}\in\mathbb{C}^{n\times n}$ is a Hermitian matrix, then the weighted 2-norm is given by $\|\mathbf{x}\|_{\mathbf{Q}}=\sqrt{\mathbf{x}^{\mathsf{H}}\mathbf{Q}\mathbf{x}}$. For $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]$, with $\mathbf{x}_{i}\in\mathbb{C}^{n}$, the operation $\operatorname{vec}(\mathbf{X})=[\mathbf{x}^{\top}_{1},\ldots,\mathbf{x}^{\top}_{n}]^{\top}$ restructures the matrix into a vector by stacking its columns.
Figure 1: Prototype experimental wafer-stage setup. The wafer chuck (1) is suspended by gravity compensators and actively controlled in six motion degrees of freedom. The sensory equipment is mounted on the metrology frame (2), while the actuators are mounted on the force frame (3). The metrology frame is suspended from the base frame (4) through an air suspension system.
@@ 2 Problem formulation
This section introduces the experimental setup and outlines the modal modeling framework for mechanical systems. Finally, the problem addressed in this paper is stated.
2.1 Experimental setup: experimental wafer-stage system
The considered experimental setup depicted in Figure 1 is a prototype wafer stage system. The stage is designed as a lightweight mechanical structure to enable fast motions with high accelerations. The system is actively controlled in six motion degrees of freedom at a sampling rate of 10 kHz, achieving accuracy in the sub-micrometer range. The stage is magnetically levitated by means of passive gravity compensators, establishing a mid-air equilibrium. The system contains 17 actuators, with 13 voice-coil actuators in the $z$-direction and two Lorentz actuators for each $x$- and $y$-directions. Furthermore, the system includes 7 sensors, 4 for the $z$-direction, 2 for the $x$-direction, and a single sensor for the $y$-direction. To facilitate the exposition, this paper only considers out-of-plane motion, that is, translation along the $z$ axis and rotations around the $x$ and $y$ axes. A schematic overview of the actuator and sensor layout in the out-of-plane direction is provided in Figure 2. The lightweight design of the stage results in high-order, and lightly-damped flexible dynamic behavior that limits system performance. Consequently, parametric models that accurately capture these dynamics are essential to enable further performance improvements.
@@ 2.2 Modal modeling of mechanical systems
This section introduces the modeling framework used to describe the wafer-stage system. The equations of motion for a spatially discretized linear-time invariant (LTI) mechanical system of order $2n$ with $n_{\mathrm{u}}$ inputs and $n_{\mathrm{y}}$ outputs are defined in terms of the generalized coordinates $\mathbf{q}(t)\in\mathbb{R}^{n}$ as follows:
@@
where the mass matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ is positive definite, the viscous damping matrix is given by $\mathbf{D}\in\mathbb{R}^{n\times n}$, the stiffness matrix $\mathbf{K}\in\mathbb{R}^{n\times n}$ is positive semi-definite, $\mathbf{F}\in\mathbb{R}^{n\times n_{\mathrm{u}}}$ is the input matrix, and $\mathbf{Q}\in\mathbb{R}^{n_{\mathrm{y}}\times n}$ is the output matrix. The coupled second-order differential equations are expressed in modal form, following standard approaches in vibration analysis Ewins2000ModalTesting ; DeKraker2004ADynamics ; Gawronski2004AdvancedStructures . Two modal formulations are considered. The first corresponds to the case of general-viscously damped systems, while the second assumes proportional damping, a special case in which the damping matrix is proportional to the mass and stiffness matrices. The choice between damping models depends on the system under consideration. While proportional damping offers a straightforward, interpretable formulation for many applications, it can be inadequate for systems involving more complex damping behavior, necessitating the use of the general form. A complete derivation of the modal forms is provided in Appendix A, and the main results are presented in the following sections.
@@ 2.2.1 General-viscously damped mechanical systems
The general case is considered first. The dynamics are partitioned into $n_{\mathrm{rbm}}$ rigid-body modes and $n_{\mathrm{flex}}$ flexible modes. A rigid-body mode represents a motion of the system involving pure translation or rotation without internal deformation and is described by second-order integrators. Since physical systems contain an infinite number of structural modes, a common approach is to estimate the dominant $n_{\mathrm{flex}}<n$ flexible modes, while approximating the influence of higher-order modes with a static gain term. This term is omitted for clarity but is straightforwardly included if required. The transfer matrix corresponding to (1), expressed in modal form with general damping, is given by
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}_{i}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{L}_{i}}{s-\lambda_{i}}+\frac{\bar{\mathbf{L}}_{i}}{s-\bar{\lambda}_{i}},$ (2)
@@
$\mathbf{R}_{i}$ $=\boldsymbol{\phi}_{l,i}\boldsymbol{\phi}_{r,i}^{\top},$ (4)
where $s\in\mathbb{C}$ is the Laplace variable, $\mathbf{R}_{i}\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\mathrm{u}}}$ and $\mathbf{L}_{i}\in\mathbb{C}^{n_{\mathrm{y}}\times n_{\mathrm{u}}}$ are rank-one modal residue matrices, $\boldsymbol{\psi}_{l,i}\in\mathbb{C}^{n_{\mathrm{y}}}$ and $\boldsymbol{\psi}_{r,i}\in\mathbb{C}^{n_{\mathrm{u}}}$ are the complex left and right mode-shape vectors of the flexible modes, $\boldsymbol{\phi}_{l,i}\in\mathbb{R}^{n_{\mathrm{y}}}$ and $\boldsymbol{\phi}_{r,i}\in\mathbb{R}^{n_{\mathrm{u}}}$ are the real left and right mode-shape vectors of the rigid-body modes, and $\lambda_{i}\in\mathbb{C}$ is the eigenvalue, given by
@@
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}_{i}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{N}_{i,1}s+\mathbf{N}_{i,0}}{s^{2}-(\lambda_{i}+\bar{\lambda}_{i})s+\lambda_{i}\bar{\lambda}_{i}},$ (6)
@@
$\boldsymbol{\rho}=\begin{bmatrix}\boldsymbol{\varrho}_{1}&\ldots&\boldsymbol{\varrho}_{n_{\mathrm{rbm}}}&\boldsymbol{\rho}_{1}&\ldots&\boldsymbol{\rho}_{n_{\mathrm{flex}}}\end{bmatrix}$. (9)
where each dynamic mode is parameterized according to
$\boldsymbol{\varrho}_{i}$ $=\begin{bmatrix}\boldsymbol{\phi}_{l,i}^{\top}&\boldsymbol{\phi}_{r,i}^{\top}\end{bmatrix},\quad\boldsymbol{\rho}_{i}=\begin{bmatrix}\Re\lambda_{i}&\Im\lambda_{i}&\Re\,\boldsymbol{\psi}_{l,i}^{\top}&\Im\,\boldsymbol{\psi}_{l,i}^{\top}&\Re\,\boldsymbol{\psi}_{r,i}^{\top}&\Im\,\boldsymbol{\psi}_{r,i}^{\top}\end{bmatrix}$. (10)
@@ 2.2.2 Proportionally damped mechanical systems
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}_{i}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{R}_{i}}{s^{2}+2\zeta_{i}\omega_{i}s+\omega_{i}^{2}},$ (11)
$\mathbf{R}_{i}$ $=\boldsymbol{\phi}_{l,i}\boldsymbol{\phi}_{r,i}^{\top},$ (12)
@@
$\boldsymbol{\rho}_{i}=\begin{bmatrix}\omega_{i}&\zeta_{i}&\boldsymbol{\phi}_{l,i}^{\top}&\boldsymbol{\phi}_{r,i}^{\top}\end{bmatrix}$. (13)
@@ 2.3 Problem statement
This paper addresses the estimation of input–output models in modal form, as defined in (9), for both proportionally damped and generally viscously damped mechanical systems using experimental data. The identification problem is structured, since the numerators are constrained to the dyadic product of mode-shape vectors. A frequency-domain identification approach is adopted, in which the objective is to estimate a model that best fits a measured Frequency Response Function (FRF) of the system. To obtain continuous-time models from FRF data, the setup outlined in (Pintelon2012SystemIdentification, , Chapter 13) is used. The resulting noisy nonparametric dataset is denoted by $\mathbf{G}(\omega_{k})\in\mathbb{C}^{n_{y}\times n_{u}}$, for $k=1,\ldots,N$, and forms the basis for subsequent parametric identification.

### v2

@@ 1 Introduction
Increasingly stringent performance requirements of next-generation industrial systems require data-driven models to integrate prior knowledge via interpretable physical principles. Embedding such principles improves model accuracy, increases interpretability, and typically reduces model complexity Soderstrom2001SystemIdentification . These characteristics are crucial not only for advanced multivariable control design methods VanDeWal2002MultivariableMotion ; VanHerpen2014ExploitingControl but also for mechanical system design, model validation, machine monitoring techniques, and simulation studies Schmidt2011TheMechatronics ; Heertjes2020ControlDevelopments ; Steinbuch2022MotionLaw ; Classens2023FromReconfiguration .
@@
Despite extensive work on modal identification, relatively few methods explicitly enforce the rank constraint inherent to modal decompositions. Traditional approaches include single-step methods, which approximate the original nonlinear least-squares problem by a linear surrogate Peeters2004TheEstimation ; Peeters2005PolyMax:Analysis . Improved estimation accuracy has been achieved through iterative multi-step methods that directly minimize the nonlinear cost function, using gradient-descent optimization Bayard1994High-orderResults , pseudo-linear regression techniques Sanathanan1961TransferPolynomials ; Blom2010MultivariableRegression or vector fitting procedures Semlyen1999RationalFitting ; Civera2021ExperimentalMethod . These approaches have been applied in a range of multi-input multi-output (MIMO) modal identification methods Verboven2005MultivariableEstimation ; Vayssettes2014Frequency-domainTests ; Reynders2012SystemComparison ; Voorhoeve2021IdentifyingStage . However, single-step methods generally lack statistical optimality guarantees, while many multi-step approaches either neglect the rank constraint or rely on singular value decompositions, which are typically effective only for systems with sufficiently spaced modes.
@@
Although many parametric modal identification techniques are available, there is currently no systematic and computationally efficient framework that explicitly enforces the rank constraints intrinsic to modal decompositions, exploits the additive structure of modal models in the model structure, and remains applicable to large-scale industrial systems with closely spaced modes. This paper introduces a novel two-stage identification strategy for estimating MIMO input-output models in modal form. In the first stage, an additive model is estimated using a refined instrumental variable method Young1980RefinedAnalysis , which is subsequently projected onto a modal form using the indirect prediction error method (IPEM) Soderstrom1991AnIdentification . The framework accommodates both proportional damping and general viscous damping models DeKraker2004ADynamics , as these formulations are commonly encountered in industrial applications Verbaan2015RobustStages . The main contributions of this paper are:
@@ 1. C1:
    A general framework is introduced for estimating continuous-time MIMO models in additive transfer-function form using a frequency-domain refined instrumental variable method.
@@ 2. C2:
    A structured identification framework is introduced that projects additive models onto the modal form, accommodating both proportional and general viscous damping model parameterizations.
@@
The remainder of this paper is organized as follows. Section 2 introduces the modal modeling framework for flexible mechanical systems, and provides a concise problem formulation. Section 3 formulates the two‐stage identification problem, with the solution procedures for the first stage provided in Section 4 and for the second stage in Section 5. Experimental validation is described in Section 6, and finally, Section 7 presents the conclusions.
Notation: Scalars, vectors, matrices and sets are written as $a$, $\mathbf{a}$, $\mathbf{A}$ and $\mathcal{A}$, respectively. The imaginary unit is denoted by $j$ and satisfies $j^{2}=-1$. For $\mathbf{z}\in\mathbb{C}^{n}$, the operation $\Re\{\mathbf{z}\}$ returns the real part, $\Im\{\mathbf{z}\}$ returns the imaginary part and $\bar{\mathbf{z}}$ returns the complex conjugate of the complex vector $\mathbf{z}$. The identity matrix of size $n$ is denoted by $\mathbf{I}_{n}$. For a matrix $\mathbf{A}$, its transpose is written as $\mathbf{A}^{\top}$, and its Hermitian (conjugate transpose) as $\mathbf{A}^{\mathsf{H}}$. If $\mathbf{x}\in\mathbb{C}^{n}$ and $\mathbf{Q}\in\mathbb{C}^{n\times n}$ is a Hermitian matrix, then the weighted 2-norm is given by $\|\mathbf{x}\|_{\mathbf{Q}}=\sqrt{\mathbf{x}^{\mathsf{H}}\mathbf{Q}\mathbf{x}}$. For $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]$, with $\mathbf{x}_{i}\in\mathbb{C}^{n}$, the operation $\operatorname{vec}(\mathbf{X})=[\mathbf{x}^{\top}_{1},\ldots,\mathbf{x}^{\top}_{n}]^{\top}$ restructures the matrix into a vector by stacking its columns.
@@ 2 Problem formulation
This section first introduces the considered modal modeling framework for describing mechanical systems and thereafter formulates the problem addressed in this paper.
@@ 2.1 Modal modeling of mechanical systems
The equations of motion for a spatially discretized linear-time invariant (LTI) mechanical system of order $2n$ with $n_{\mathrm{u}}$ inputs and $n_{\mathrm{y}}$ outputs are defined in terms of the generalized coordinates $\mathbf{q}(t)\in\mathbb{R}^{n}$ as follows
@@
with $\mathbf{y}(t)\in\mathbb{R}^{n_{\mathrm{y}}}$ the system outputs, $\mathbf{u}(t)\in\mathbb{R}^{n_{\mathrm{u}}}$ the system inputs, the mass matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ is positive definite, the viscous damping matrix is given by $\mathbf{D}\in\mathbb{R}^{n\times n}$, the stiffness matrix $\mathbf{K}\in\mathbb{R}^{n\times n}$ is positive semi-definite, $\mathbf{F}\in\mathbb{R}^{n\times n_{\mathrm{u}}}$ is the input matrix, and $\mathbf{Q}\in\mathbb{R}^{n_{\mathrm{y}}\times n}$ is the output matrix. The coupled second-order differential equations are expressed in modal form, following standard approaches in vibration analysis Ewins2000ModalTesting ; DeKraker2004ADynamics ; Gawronski2004AdvancedStructures . Two modal formulations are considered. The first corresponds to generally viscously damped systems, while the second assumes proportional damping, a special case in which the damping matrix is proportional to the mass and stiffness matrices, resulting in mode-shape vectors that are real-valued and coincide with those of the undamped system Gawronski2004AdvancedStructures . The choice between damping models depends on the system under consideration. While proportional damping provides a simpler and more interpretable model description due to the real-valued mode-shape vectors, it can be inadequate for systems involving more complex damping behavior, necessitating the use of the general form (see, e.g., Verbaan2015RobustStages ). A complete derivation of the modal forms is provided in Appendix A, and the main results are presented in the following subsections.
@@ 2.1.1 General-viscously damped mechanical systems
The general case is considered first. The dynamics are partitioned into $n_{\mathrm{rbm}}$ rigid-body modes and $n_{\mathrm{flex}}$ flexible modes, where $n=n_{\mathrm{rbm}}+n_{\mathrm{flex}}$. A rigid-body mode represents a motion of the system involving pure translation or rotation without internal deformation and is described by a second-order model with two poles at $s=0$. While physical systems contain infinitely many dynamic modes, only the dominant modes within the excited and measured frequency range are estimated explicitly. Modes at substantially higher frequencies typically have only a minor effect in this range, and their combined contribution can be approximated by a residual static gain Gawronski2004AdvancedStructures . This term is omitted for notational clarity but is straightforwardly included if required. The transfer matrix corresponding to (1), expressed in modal form with general damping, is given by
$\mathbf{P}_{m}(s)$ $=\sum_{j=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}^{\mathrm{rbm}}_{j}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{L}_{i}}{s-\lambda_{i}}+\frac{\bar{\mathbf{L}}_{i}}{s-\bar{\lambda}_{i}},$ (2)
@@
$\mathbf{R}^{\mathrm{rbm}}_{j}$ $=\boldsymbol{\phi}_{l,j}\boldsymbol{\phi}_{r,j}^{\top},$ (4)
where $s\in\mathbb{C}$ is the Laplace variable, $\mathbf{R}^{\mathrm{rbm}}_{j}\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\mathrm{u}}}$ and $\mathbf{L}_{i}\in\mathbb{C}^{n_{\mathrm{y}}\times n_{\mathrm{u}}}$ are rank-one modal residue matrices, $\boldsymbol{\psi}_{l,i}\in\mathbb{C}^{n_{\mathrm{y}}}$ and $\boldsymbol{\psi}_{r,i}\in\mathbb{C}^{n_{\mathrm{u}}}$ are the complex left and right mode-shape vectors of the flexible modes, $\boldsymbol{\phi}_{l,j}\in\mathbb{R}^{n_{\mathrm{y}}}$ and $\boldsymbol{\phi}_{r,j}\in\mathbb{R}^{n_{\mathrm{u}}}$ are the real left and right mode-shape vectors of the rigid-body modes, and $\lambda_{i}\in\mathbb{C}$ is the eigenvalue, given by
@@
$\mathbf{P}_{m}(s)$ $=\sum_{j=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}^{\mathrm{rbm}}_{j}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{N}_{i,1}s+\mathbf{N}_{i,0}}{s^{2}-(\lambda_{i}+\bar{\lambda}_{i})s+\lambda_{i}\bar{\lambda}_{i}},$ (6)
@@
$\boldsymbol{\rho}=\begin{bmatrix}\boldsymbol{\varrho}^{\top}_{1}&\ldots&\boldsymbol{\varrho}^{\top}_{n_{\mathrm{rbm}}}&\boldsymbol{\rho}^{\top}_{1}&\ldots&\boldsymbol{\rho}^{\top}_{n_{\mathrm{flex}}}\end{bmatrix}^{\top}$. (9)
where the vector $\boldsymbol{\varrho_{j}}$ denotes the parameter vector of the $j$th rigid-body mode, and $\boldsymbol{\rho}_{i}$ denotes the parameter vector of the $i$th flexible dynamic mode, which are both given by
$\boldsymbol{\varrho}_{j}$ $=\begin{bmatrix}\boldsymbol{\phi}_{l,j}^{\top}&\boldsymbol{\phi}_{r,j}^{\top}\end{bmatrix}^{\top},\quad\boldsymbol{\rho}_{i}=\begin{bmatrix}\Re\lambda_{i}&\Im\lambda_{i}&\Re\,\boldsymbol{\psi}_{l,i}^{\top}&\Im\,\boldsymbol{\psi}_{l,i}^{\top}&\Re\,\boldsymbol{\psi}_{r,i}^{\top}&\Im\,\boldsymbol{\psi}_{r,i}^{\top}\end{bmatrix}^{\top}$. (10)
@@ 2.1.2 Proportionally damped mechanical systems
$\mathbf{P}_{m}(s)$ $=\sum_{j=1}^{n_{\mathrm{rbm}}}\frac{\mathbf{R}^{\mathrm{rbm}}_{j}}{s^{2}}+\sum_{i=1}^{n_{\mathrm{flex}}}\frac{\mathbf{R}^{\mathrm{flex}}_{i}}{s^{2}+2\zeta_{i}\omega_{i}s+\omega_{i}^{2}},$ (11)
$\mathbf{R}^{\mathrm{flex}}_{i}$ $=\boldsymbol{\phi}_{l,i}\boldsymbol{\phi}_{r,i}^{\top},$ (12)
@@
$\boldsymbol{\rho}_{i}=\begin{bmatrix}\omega_{i}&\zeta_{i}&\boldsymbol{\phi}_{l,i}^{\top}&\boldsymbol{\phi}_{r,i}^{\top}\end{bmatrix}^{\top}$. (13)
@@ 2.2 Problem statement
This paper addresses the estimation of input–output models in modal form, as defined in (9), for both proportionally damped and generally viscously damped mechanical systems using experimental data. The identification problem is structured, since the numerators are constrained to the dyadic product of the mode-shape vectors. A frequency-domain identification approach is adopted, in which the objective is to estimate a model that best fits a measured Frequency Response Function (FRF) of the system. To obtain continuous-time models from FRF data, the setup outlined in (Pintelon2012SystemIdentification, , Chapter 13) is used. The noisy FRF dataset is denoted by $\{\mathbf{G}(\omega_{k})\}_{k=1}^{N}$, with $N$ the number of frequency points and $\mathbf{G}(\omega_{k})\in\mathbb{C}^{n_{y}\times n_{u}}$ the response at frequency $k$, and forms the basis for subsequent parametric identification.